### v1

# Unimodular polytopes and a new Heller-type bound on totally unimodular matrices via Seymour’s decomposition theorem

Abstract.
We prove a sharp upper bound on the number of columns of a totally unimodular matrix with column sums $1$ which improves upon Heller’s bound by a factor of $2$. The proof uses Seymour’s decomposition theorem. Unimodular polytopes are lattice polytopes such that the vertices of every full-dimensional subsimplex form an affine lattice basis. This is a subclass of 0/1-polytopes and contains for instance edge polytopes of bipartite graphs. Our main result implies a sharp upper bound on the number of vertices of unimodular polytopes.
1. A Heller-type bound on polytopal TU-matrices
Totally unimodular matrices are objects of fundamental importance in optimization and matroid theory [15]. Let us recall their definition.
Definition 1.1.
An $m\times n$-matrix $A$ with entries in $\mathbb{Z}$ is called totally unimodular or TU-matrix if every minor is in $\{-1,0,1\}$.
Note that particularly every entry of a TU-matrix is in $-1,0,-1$. Moreover, every submatrix of a TU-matrix is again TU.
Heller [8] showed that any TU-matrix with $m$ rows such that all columns are pairwise distinct has at most $m^{2}+m+1$ columns. To see the similarity to our main result below let us write this in the following equivalent form.
Theorem 1.2 (Heller’s bound).
Let $M$ be a TU matrix with $m$ rows and $n$ columns such that all columns are pairwise distinct. If there exists $f\in(\mathbb{Z}^{m})^{*}$ such that $f$ evaluates to a positive number on every column of $M$ (e.g., all column sums are positive), then $n\leq\frac{(m+1)m}{2}$.
To see the equivalence to above original formulation of Heller, for the one direction consider the extended matrix consisting of $M$, $-M$, and the zero column, while for the reverse direction note that one can always separate a finite number of non-zero vectors by an integral functional into two halfspaces.
As the special case of our interest, let us introduce the following terminology.
Definition 1.3.
An $m\times n$-matrix TU-matrix $M$ is called polytopal if there exists $f\in(\mathbb{Z}^{m})^{*}$ such that $f$ evaluates to $1$ on every column of $M$.
Because of the TU-property this is equivalent to all columns lying in a common affine hyperplane that does not contain the origin. An important special case is that all column sums are $1$. Here is our main result.
Theorem 1.4.
Let $M$ be a TU matrix with $m$ rows and $n$ columns such that all columns are pairwise distinct. If $M$ is polytopal (e.g., all column sums are $1$), then $n$ is bounded from above by $10$ if $m=5$ and by $\lfloor\frac{(m+1)^{2}}{4}\rfloor$ if $m\not=5$.
We remark that this sharpens in this polytopal situation Heller’s bound (roughly) by a factor of $2$.
Example 1.5.
The bound in Theorem 1.4 is sharp for all $m$ as the following examples show. For $m=5$ consider the following TU-matrix:
$\begin{pmatrix}1&0&0&0&0&1&0&0&1&-1\\ 0&1&0&0&0&-1&1&0&0&1\\ 0&0&1&0&0&1&-1&1&0&0\\ 0&0&0&1&0&0&1&-1&1&0\\ 0&0&0&0&1&0&0&1&-1&1\end{pmatrix}$
[…]
As the exceptional case for $m=5$ already hints at, the proof uses the celebrated decomposition theorem of Seymour.
Our motivation for the study of polytopal TU-matrices comes from their direct relation to unimodular polytopes, see Proposition 2.9 below. In the next section we will discuss this interesting class of lattice polytopes and use Theorem 1.4 to deduce a sharp upper bound on their number of vertices in Theorem 2.7. The proof of Theorem 1.4 will occupy Section 3. We prove it first in the setting of network matrices and transposes of network matrices, while we rely on Seymour’s decomposition theorem to prove it in the general case.
Remark 1.6.
We believe that our bound in Theorem 1.4 also holds for TU-matrices such that all columns are pairwise distinct and have positive odd column sums. In fact, the proof of Theorem 1.4 could be immediately extended to this case, the only missing part is the computational task to verify that it holds in the case of at most $6$ rows (which is true for column sums equal to $1$ by Remark 3.13). This will be done in a future version of this paper.
Acknowledgment
We thank Gennadiy Averkov, Volker Kaibel and Matthias Schymura for their interest and remarks. The author is funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – 539867500 as part of the research priority program Combinatorial Synergies.
2. Unimodular polytopes
In this section, we will apply Theorem 1.4 to the setting of unimodular polytopes.
2.1. Definition
Let $P\subset\mathbb{R}^{d}$ be a lattice polytope, i.e., $P$ is the convex hull of points in the lattice $\mathbb{Z}^{d}$. Two lattice polytopes are isomorphic if there is an affine-linear automorphism of $\mathbb{Z}^{d}$ mapping the vertices bijectively to each other.
A lattice simplex $P$ of dimension $d$ is called unimodular if its vertices form an affine lattice basis of $\mathbb{Z}^{d}$. Equivalently, $P\cong\Delta_{d}:=\operatorname{conv}(0,e_{1},\ldots,e_{d})$. Note that a unimodular simplex contains only its vertices as lattice points. The definition of unimodular simplices can be extended to lattice polytopes [4, Sect. 6.2].
Definition 2.1.
$P$ is called unimodular polytope if every full-dimensional subsimplex of the vertices of $P$ is unimodular.
Note that the only lattice points in a unimodular polytope are its vertices.
Remark 2.2.
Let us give some equivalent conditions that are well-known or straightforward to check. For this, let us recall that a lattice triangulation of $P$ is called unimodular if every full-dimensional cell in the triangulation is a unimodular simplex.
  1. (1)
    $P$ is unimodular
  2. (2)
    every lattice subpolytope of $P$ is unimodular
  3. (3)
    every lattice triangulation of $P$ is unimodular
  4. (4)
    every regular lattice triangulation of $P$ is unimodular
2.2. Classification results and examples
Let us note the following observation.
Proposition 2.3.
$P$ is unimodular and simple if and only if $P$ is a unimodular simplex or isomorphic to a product of two unimodular simplices.
Proof.
The ’if’-direction has been proven in [4, Proposition 6.2.1]. On the other hand, if $P$ is unimodular and simple it is also smooth (i.e., the convex hull of every vertex and the closest lattice points on the incident edges is a unimodular simplex). It has been shown in [9, Theorem 3.32] that every smooth polytope that has only its vertices as lattice points is a product of unimodular simplices. Now, we just have to observe that every product of more than two unimodular simplices contains $[0,1]^{3}$ as a face, which is not unimodular. ∎
Example 2.4.
Consider an undirected bipartite graph $G$ with parts $A$ and $B$. Then the edge polytope of $G$ is defined as the convex hulls of the vectors $e_{i}+e_{j}$ associated to edges with vertices $i$ and $j$. For the complete bipartite graph $K_{a,b}$, we have $P(K_{a,b})\cong\Delta_{a-1}\times\Delta_{b-1}$, hence, any edge polytope of a bipartite graph is unimodular.
Unimodular polytopes form a special subclass of compressed polytopes (i.e., every pulling triangulation is unimodular) [4, Sect. 9.3]. In particular, they are 0/1-polytopes, i.e., they are isomorphic to lattice subpolytopes of $[0,1]^{\dim(P)}$. Note that $[0,1]^{d}$ is compressed but not unimodular if $d\geq 3$.
0/1-polytopes are classified up to dimension $5$ [1] and available in the database PolyDB [14]. Using Polymake [6] and OSCAR [13, 3] we get the following complete classification of unimodular polytopes.
Proposition 2.5.
For each dimension $d$ there are the following number of isomorphism classes of unimodular polytopes of dimension $d$: 1 2 3 4 5 1 2 4 13 38
2.3. The maximal number of lattice points/vertices
As the maybe most special non-trivial class of lattice polytopes, unimodular polytopes should be well studied, however, surprisingly many elementary questions are still open. Here, we are interested in determining the maximal number of vertices (same as the number of lattice points). One may guess that the maximum is always attained by the product of two unimodular simplices, however, this is wrong in dimension $d=4$, where the number of vertices of $\Delta_{2}\times\Delta_{2}$ is $9$.
Example 2.6.
The convex hull of the columns of the following $0/1$-matrix is a $4$-dimensional unimodular polytope with $10$ vertices: $\begin{pmatrix}0&0&0&0&1&1&1&1&1&1\\ \\ 0&1&1&1&0&0&0&1&1&1\\ \\ 1&0&1&1&0&1&1&0&0&1\\ \\ 1&1&0&1&1&0&1&0&1&0\end{pmatrix}$
As we will explain below, applying our main result on polytopal TU-matrices implies a sharp bound on the number of vertices of unimodular polytopes that shows that this exceptional behavior only happens in dimension $4$.
Theorem 2.7.
The number of vertices of a $d$-dimensional unimodular polytope $P$ is bounded from above by $10$ if $d=4$ and by $\lfloor\frac{(d+2)^{2}}{4}\rfloor$ otherwise. The bound is attained by Example 2.6 if $d=4$, by $\Delta_{\frac{d}{2}}\times\Delta_{\frac{d}{2}}$ if $d\not=4$ even, and by $\Delta_{\frac{d+1}{2}}\times\Delta_{\frac{d-1}{2}}$ if $d$ is odd.
Remark 2.8.
It seems a very challenging open question to find similar sharp results for $\Delta$-modular polytopes: lattice polytopes where $\Delta$ is the maximal normalized volume of any full-dimensional lattice subsimplex. This question is motivated by recent papers [7, 11, 2, 12] on the size of integer programs with bounded subdeterminants.
2.4. The relation to TU-matrices
Theorem 2.7 follows immediately from Theorem 1.4. Let us explain this direct connection between unimodular polytopes and TU-matrices in more detail. We abbreviate $\operatorname{conv}(M)$ for the convex hull of the columns of a matrix $M$.
Proposition 2.9.
Let $M$ be a polytopal $(d+1)\times n$ TU-matrix of rank $d+1$. Then $\operatorname{conv}(M)\subset\mathbb{R}^{d+1}$ is a unimodular polytope of dimension $d$ with $n$ vertices.
On the other hand, every unimodular polytope of dimension $d$ is isomorphic to $\operatorname{conv}(I_{d+1}|B)$ where $B$ and $(I_{d+1}|B)$ are TU-matrices with column sums $1$.
Proof.
All $d$-dimensional subsimplices of the columns of $M$ are unimodular as their vertices form a lattice basis of $\mathbb{Z}^{d+1}$ by the TU-property. Hence, $\operatorname{conv}(M)$ is a unimodular polytope of dimension $d$.
On the other hand, let $P$ be a unimodular polytope of dimension $d$. We may assume that $P\subset\mathbb{R}^{d}\times\{1\}$ and that the first $d+1$ vertices of $P$ form a lattice basis of $\mathbb{Z}^{d+1}$. Let $M$ be the matrix with the vertices of $P$ as columns. Multiplying from the left with the inverse of the matrix of the first $d+1$ columns yields a matrix $M^{\prime}=(I_{d+1}|B)$ with $P\cong\operatorname{conv}(M^{\prime})$ where $M^{\prime}$ necessarily has column sums $1$. As $P$ is unimodular, all $(d+1)$-minors of $M^{\prime}$ are in $\{-1,0,1\}$. Note that the $(d+1)$-minors of $M^{\prime}$ with $k$ columns in $I_{d+1}$ equal $\pm$ the $(d+1-k)$-minors of $B$. Hence, $B$ is a TU-matrix and also $(I_{d+1}|B)$ is a TU-matrix. ∎
Hence, we see that Theorem 1.4 implies Theorem 2.7 with $m=d+1$.
3. The proof of the Heller-type bound on TU-matrices
The remainder of this paper contains the proof of Theorem 1.4 using Seymour’s decomposition theorem. First, we will look at network matrices and their transposes, before dealing with the inductive step using so-called $k$-sums.
3.1. Network matrices
We say $M$ is an $A_{0}\times A$-network matrix [15] (or tree matrix [17]) if it is associated to the set of arcs $A_{0}$ of a directed tree $T$ with vertex set $V$ and to the set of arcs $A$ of a digraph $D$ on $V$ in the following way. Let $e$ be an arc of $D$ with vertices $s\rightarrow t$, and let $H$ be the unique path in the underlying undirected graph of $T$ from $s$ to $t$. Then the column of $M$ corresponding to $e\in A$ contains at the row position corresponding to an arc $e^{\prime}$ of $T$ the entry $1$ if $e^{\prime}$ is passed with the same orientation in $H$, $-1$ if it is passed with the reverse orientation in $H$, and 00 if this edge is not used in $H$.
Proposition 3.1.
Let $M$ be an $A_{0}\times A$-network matrix such that all columns are pairwise different and have positive odd column sums. Then $|A|$ is bounded from above by $\frac{(|A_{0}|+1)^{2}}{4}$. Equality holds if and only if $|A_{0}|$ is odd and the underlying digraph $D$ is a complete bipartite graph with two parts of the same size $\frac{|A_{0}|+1}{2}$.
Proof.
We use notation and results from Schrijver’s book [15] (see proof of equation (34) on page 277). Let $N$ be the $V\times A$-incidence matrix of $D$ and $L$ be the $V\times A_{0}$-incidence matrix of $T$. Moreover, let $\tilde{N}$ and $\tilde{L}$ be given by deleting the last row in $N$, respectively $L$. Note that $LM=N$, $\tilde{L}M=\tilde{N}$, and $\tilde{L}$ is invertible. Let $\mathbf{1}:=(1,\ldots,1)\in\mathbb{Z}^{A_{0}}$ and $z=(z_{1},\ldots,z_{A})\in\mathbb{Z}^{A}$ such that $\mathbf{1}M=z^{t}$. Thus, $\mathbf{1}\tilde{L}^{-1}\tilde{N}=z^{t}$. Define $u\in\mathbb{Z}^{V}$ as $\mathbf{1}\tilde{L}^{-1}$ extended by 00 in the last coordinate. Then we see $u^{t}N=z^{t}$.
Let $G$ be the undirected graph underlying $D$. We will show that $G$ is bipartite by observing that all cycles must be even. For this, let $v_{1},\ldots,v_{r}$ be vertices in $V$ forming a cycle in $G$ and assume that $r$ is odd. We set $v_{r+1}:=v_{1}$. Denoting the successive corresponding arcs in $A$ by $a_{1},\ldots,a_{r}$, we get from $u^{t}N=z^{t}$ that $u_{v_{i+1}}=u_{v_{i}}\pm z_{a_{i}}$ for $i=1,\ldots,r$. Hence, $u_{v_{r+1}}=u_{v_{1}}\pm z_{a_{1}}\pm\cdots\pm z_{a_{r}}$, so ±$\pm z_{a_{1}}\pm\cdots\pm z_{a_{r}}=0$. As each summand is odd and $r$ is odd, we derive at a contradiction.
It is well-known that a bipartite graph $G$ with $|V|$ vertices can have at most $\frac{|V|^{2}}{4}$ edges, with equality if and only if $|V|$ is even and $G$ is a complete bipartite graph with two parts of size $\frac{|V|}{2}$. Now, use $|V|=|A_{0}|+1$. ∎
3.2. Transposes of network matrices
Theorem 3.2.
Let $M$ be an $A_{0}\times A$-network matrix. Then the set of pairwise different rows of $M$ with positive odd row sums is bounded from above by $\frac{(|A|+1)^{2}}{4}$ if $|A|$ is odd, respectively, $\frac{|A|^{2}}{4}$ if $|A|$ is even. Moreover, equality implies that $|A|\leq 6$.
For the proof we reduce the problem to one about undirected graphs. Let us recall that $[m]:=\{1,\ldots,m\}$.
Proposition 3.3.
Let $T$ be an (undirected) tree that contains $m$ (undirected) paths $P_{1},\ldots,P_{m}$. Set
$I_{e}:=\{j\in[m]\;:\;e\in E(P_{j})\}.$
Then for each $k\in[m]$ with $k\not=2$ we have
|$|\{I_{e}\;:\;e\in E(T),|I_{e}|=k\}|\leq m+1-k.$
Moreover, there cannot be simultaneously equality for $k=5$ and $k=7$.
The proof follows from considering a more special situation.
Lemma 3.4.
In the situation of Proposition 3.3, let each edge $e$ of $T$ be contained in precisely $k$ of these $m$ paths. If $k\not=2$, then we have
|$|\{I_{e}\;:\;e\in E(T)\}|\leq m+1-k.$
Moreover, equality for $k\geq 5$ implies that $T$ is a path, each of the $m$ paths contains at least one edge, and $|I_{e}\cap I_{e-1}|\in\{k-1,k\}$ for every pair $e,e^{\prime}\in E(T)$, $e\not=e^{\prime}$ with $V(e)\cap V(e^{\prime})\not=\emptyset$.
Proof.
If $k=1$, then trivially |$|\{I_{e}\;:\;e\in E(T)\}|=|[m]|=m=m+1-k$. Hence, let $k\geq 3$. We prove this by induction on the number of vertices of $T$. If $V(T)\leq 2$, the statement is immediate. So, let $V(T)\geq 3$. There exists an internal vertex $v\in V(T)$ that we consider as a root of $T$. Let $r\geq 2$ be the number of incident edges, i.e., children. Denote by $T_{1},\ldots,T_{r}$ the $r$ subtrees of $T$ originating at $v$. Let us observe that each path can only use edges from at most two of these subtrees. Let us denote for $i\in[r]$ by $P_{i,i}$ the set of paths that use only edges from $T_{i}$. For $i,j\in[r]$ with $i\not=j$ let $P_{i,j}$ denote the set of paths that use edges in $T_{i}$ as well as in $T_{j}$. Applying the induction hypothesis to each $T_{i}$, we get
|$|\{I_{e}\;:\;e\in E(T_{i})\}|\leq\left(\sum_{j=1}^{r}|P_{i,j}|\right)+1-k.$
Summing this up, we get
|$|\{I_{e}\;:\;e\in E(T)\}|\leq\left(\sum_{i,j}|P_{i,j}|\right)+r(1-k).$
Note that $m=\left(\sum_{i=1}^{r}|P_{i,i}|\right)+(\sum_{i<j}|P_{i,j}|)$. Putting this into the right hand side we get $m+\left(\sum_{i<j}|P_{i,j}|+r(1-k)\right)$. Hence, we see that it suffices to prove
$\sum_{i<j}|P_{i,j}|\leq(r-1)(k-1).$
For each $i\in[r]$ let $e_{i}\in V(T_{i})$ be the unique edge incident to $v$. Then $e_{i}$ is contained in at least $\sum_{j\not=i}|P_{i,j}|$ many paths. Hence, $\sum_{j\not=i}|P_{i,j}|\leq k$ for each $i\in[r]$. Adding this up and dividing by $2$, we get
$\sum_{i<j}|P_{i,j}|\leq\frac{rk}{2}.$
If $r\geq 3$ and $k\geq 4$ (or, $r\geq 4$ and $k=3$), then $r\geq 2\frac{k-1}{k-2}$, which implies $\sum_{i<j}|P_{i,j}|\leq\frac{rk}{2}\leq(r-1)(k-1)$ as desired.
If $r=3=k$, then we would like to show that the sum of $|P_{1,2}|$,$|P_{1,3}|$,$|P_{2,3}|$ is at most $(r-1)(k-1)=4$. However, this follows simply from the fact that the sum of two of these summands cannot exceed $k=3$.
The last case to consider is that there is no $v$ with $r\geq 3$, i.e., all internal vertices of $T$ have degree $2$. This implies that $T$ is a path. Except for the trivial case |$|\{I_{e}\;:\;e\in E(T)\}|=1$, we can now choose $v$ with $r=2$ such that $I_{e_{1}}\not=I_{e_{2}}$. In particular, $|P_{1,2}|<k$. Thus, $|P_{1,2}|\leq k-1=(r-1)(k-1)$, as desired.
Finally, let us assume that we have |$|\{I_{e}\;:\;e\in E(T)\}|=m+1-k$ and $k\geq 5$. Equality in above inequalities implies that $(r,k)\in\{(4,3),(3,4),(3,3)\}$ or $T$ is a path with the desired property. ∎
Proof of Proposition 3.3.
Fix some $k\in[m]$, $k\not=2$. We proceed by induction on the number of vertices of $T$. The statement is easily checked for $|V(T)|\leq 2$. Let $|V(T)|\geq 3$. By Lemma 3.4 we may assume that there exists $e\in E(T)$ such that $|I_{e}|\not=k$. By compressing $e$ in $T$ and $P_{1},\ldots,P_{m}$, the desired bound follows from applying the induction hypothesis to the resulting tree with one less vertex.
Let us call an edge $e\in T$ an $j$-edge if $|I_{e}|=j$. We assume that we have equality in the proven bound for $k=5$ and $k=7$. By Lemma 3.4, compressing all $j$-edges with $j\not\in\{5,7\}$ we get a subtree $T^{\prime}$ of $T$ consisting solely of $5$-edges and $7$-edges, such that compressing all $5$-edges yields a path of $7$-edges and compressing all $7$-edges a path of $5$-edges. Clearly, there exists a $5$-edge $e$ incident to a $7$-edge $e^{\prime}$. Let $J:=I_{e^{\prime}}\setminus I_{e}$, so $|J|\geq 2$. As each path $P_{j}$ ($j\in J$) contains at least one $5$-edge, we see that there exists a path in $T^{\prime}$ starting with $e$, followed possibly by some $7$-edges that eventually ends with a $5$-edge $e^{\prime\prime}$ with $J\subseteq I_{e^{\prime\prime}}$. Hence, $|I_{e^{\prime\prime}}\setminus I_{e}|\geq 2$, a contradiction to $|I_{e}\cap I_{e^{\prime\prime}}|\geq 4$. ∎
Proof of Theorem 3.2.
Let $m:=|A|$. We consider the underlying undirected graph of $T$ and the $m$ paths $P_{1},\ldots,P_{m}$ on $T$ associated to the $m$ edges of $A$. We define $\bar{M}$ as the $|E(T)|\times m$-matrix with entries in $\{0,1\}$ given by reducing any entry of $M$ modulo $2$. Note that the edge $e$ of $T$ lies on $P_{j}$ for $j\in[m]$ if and only if the corresponding entry in $\bar{M}$ is equal to $1$. Hence, $I_{e}$ equals the ‘support set’ of the corresponding row in $\bar{M}$. Let us also observe that, as each row vector of $M$ has only entries in $-1,0,+1$, having odd row sum in $M$ is equivalent to odd row sum in $\bar{M}$. Now, Proposition 3.3 implies that the number of pairwise different rows of $\bar{M}$ with odd row sum is bounded from above by
$\sum_{k\in[m],\,k\text{ odd}}m+1-k,$
which equals $(\frac{m+1}{2})^{2}$ if $m$ is odd, respectively, $(\frac{m}{2})^{2}$ if $m$ is even.
For the desired bound it suffices to show that distinct rows in $M$ with positive row sum stay distinct in $\bar{M}$. This implies the result. Assume that there are two arcs $e$ and $e^{\prime}$ in $A_{0}$ whose corresponding rows in $M$ have positive rows sum that have the same row in $\bar{M}$, i.e., $I_{e}=I_{e^{\prime}}$. If there would be two paths $P_{i}$ and $P_{j}$ in $T$ (for $i,j\in I_{e}=I_{e^{\prime}},\,i\not=j$) such that $M_{e,i}=M_{e,j}$ but $M_{e^{\prime},i}\not=M_{e^{\prime},j}$, then this would lead to a contradiction (e.g., by violating the necessary $2\times 2$-criterion for total unimodularity). Hence, the $e$-row of $M$ equals $\pm$ the $e^{\prime}$-row of $M$. As both rows have positive row sum, we see that they must be equal.
For the equality case we note that equality for $m\geq 7$ in Theorem 3.2 would imply equality in Proposition 3.3 for $k=5$ and $k=7$ which is not possible. ∎
Remark 3.5.
It would be interesting to see whether Proposition 3.3 can be more immediately derived from results on circuit-trees, see e.g. [5].
3.3. On $k$-sums of TU-matrices
Let us recall the $k$-sum construction, where we follow mainly the notation given in Seymour’s survey article [17]. By convention, vectors are considered as column vectors.
In the following definitions, we assume that $A$ is an $m_{1}\times n_{1}$-matrix and $B$ is an $m_{2}\times n_{2}$-matrix.
Definition 3.6.
We say
$\begin{pmatrix}A&0\\ 0&B\end{pmatrix}$
is a $1$-sum of TU-matrices $A$ and $B$.
Definition 3.7.
We say
$\begin{pmatrix}A&uv^{t}\\ 0&B\end{pmatrix}$
is a $2$-sum of TU-matrices $(A|u)$ and $\begin{pmatrix}v^{t}\\ B\end{pmatrix}$.
Note that adding a standard basis vector to a TU-matrix as a new column is a $2$-sum construction (e.g., choose $A=1=u$ in order to add the first standard basis vector).
Recall that the size of an $m\times n$-matrix is defined as $m+n$. The additional assumption in the two following definitions correspond precisely to the condition that the size of each ‘factor’ TU-matrix in a $k$-sum must be strictly smaller than the size of the whole TU-matrix.
Definition 3.8.
If the sizes of $A$ and $B$ are at least 4, we say
$\begin{pmatrix}A&C\\ 0&B\end{pmatrix}$
is a $3$-sum of TU-matrices $(A|u_{1}|u_{2}|u_{3})$ and $\begin{pmatrix}v^{t}_{1}\\ v^{t}_{2}\\ v^{t}_{3}\\ B\end{pmatrix}$, where $u_{1}+u_{2}+u_{3}=0$ and $v_{1}+v_{2}+v_{3}=0$, and every column of $C$ is one of ±$\pm u_{1},\pm u_{2},\pm u_{3},0$ and every row of $C$ is one of ±$\pm v^{t}_{1},\pm v^{t}_{2},\pm v^{t}_{3},0$.
Definition 3.9.
If the sizes of $A$ and $B$ are at least 4, we say
$\begin{pmatrix}A&u^{\prime}v^{t}\\ v^{\prime}u^{t}&B\end{pmatrix}$
is a $4$-sum of TU-matrices $\begin{pmatrix}A&u^{\prime}&u^{\prime}\\ u^{t}&0&x\end{pmatrix}$ and $\begin{pmatrix}v^{t}&0&x\\ B&v^{\prime}&v^{\prime}\end{pmatrix}$ with $x=\pm 1$.
Let us introduce some more terminology.
Definition 3.10.
Let $M$ be an $m\times n$-matrix TU-matrix. We say $M$ is $w$-valued for $w\in\mathbb{Z}^{n}$ if there exists $f\in(\mathbb{Z}^{m})^{*}\setminus\{0\}$ such that $f$ evaluates on the $j$th column of $M$ to $w_{j}$ for any $j\in[n]$.
Lemma 3.11.
Let $M$ be a TU $m\times n$-matrix that is $w$-valued with functional $f$ and has distinct columns such that $M$ is a $k$-sum (for some $k\in[4]$) in above notation. Let $w=(w_{1},w_{2})$ for $w_{1}\in\mathbb{Z}^{n_{1}}$ and $w_{2}\in\mathbb{Z}^{n_{2}}$.
  1. (1)
    If $k=2$, then $\begin{pmatrix}v^{t}\\ B\end{pmatrix}$ has distinct columns, and there is a functional $f^{\prime}\in(\mathbb{Z}^{m_{2}+1})^{*}$ such that $\begin{pmatrix}v^{t}\\ B\end{pmatrix}$ is $w_{2}$-valued, and $f^{\prime}(e_{j})$ for $j=2,\ldots,m_{2}+1$ equals $f(e_{m_{1}+j-1})$.
  2. (2)
    If $k=3$ and ±$\pm v^{t}_{1}$ and ±$\pm v^{t}_{2}$ appear as rows of $C$, then $\begin{pmatrix}v^{t}_{1}\\ v^{t}_{2}\\ B\end{pmatrix}$ has distinct columns, and there is a functional $f^{\prime}\in(\mathbb{Z}^{m_{2}+2})^{*}$ such that $\begin{pmatrix}v^{t}_{1}\\ v^{t}_{2}\\ B\end{pmatrix}$ is $w_{2}$-valued, and $f^{\prime}(e_{j})$ for $j=3,\ldots,m_{2}+2$ equals $f(e_{m_{1}+j-2})$.
  3. (3)
    If $k=4$, then $\begin{pmatrix}A\\ u^{t}\end{pmatrix}$ and $\begin{pmatrix}v^{t}\\ B\end{pmatrix}$ have each distinct columns, there is a functional $f^{\prime}\in(\mathbb{Z}^{m_{1}+1})^{*}$ such that $\begin{pmatrix}A\\ u^{t}\end{pmatrix}$ is $w_{1}$-valued and $f^{\prime}(e_{j})$ for $j=1,\ldots,m_{1}$ equals $f(e_{j})$, and there is a functional $f^{\prime\prime}\in(\mathbb{Z}^{m_{2}+1})^{*}$ such that $\begin{pmatrix}v^{t}\\ B\end{pmatrix}$ is $w_{2}$-valued and $f^{\prime\prime}(e_{j})$ for $j=2,\ldots,m_{2}+1$ equals $f(e_{m_{1}+j-1})$.
Proof.
We use $f$ from Definition 3.10.
Let $k=2$. Assume that two columns $j_{1},j_{2}$ of $\begin{pmatrix}v^{t}\\ B\end{pmatrix}$ would be equal. Then also the $(n_{1}+j_{1})$-th and the $(n_{1}+j_{2})$-th column of $M$ would be equal to each other, a contradiction. , and $f=(f_{1},f_{2})$ with $f_{1}\in(\mathbb{Z}^{m_{1}})^{*}$ and $f_{2}\in(\mathbb{Z}^{m_{2}})^{*}$. We observe that for $j\in[n_{2}]$, the $f$-value of the $(n_{1}+j)$-th column of $M$ equals the sum of $f_{1}(u)\cdot v_{j}$ and the $f_{2}$-value of the $j$-th column of $B$. For $f^{\prime}=(f_{1}(u),f_{2})\in(\mathbb{Z}^{m_{2}+1})^{*}$ the statements follow.
Let $k=3$. Assume that two columns $j_{1},j_{2}$ of $\begin{pmatrix}v^{t}_{1}\\ v^{t}_{2}\\ B\end{pmatrix}$ would be equal. Then $v_{1}+v_{2}+v_{3}=0$ implies that the $(n_{1}+j_{1})$-th and the $(n_{1}+j_{2})$-th column of $M$ would be equal to each other, a contradiction. Here, we use the coordinate-wise notation $f=(f_{1},\ldots,f_{m})$. Let us define $f^{\prime}$ as follows. For $k\in[3]$, we define $I_{k}\subseteq[m_{1}]$ as the set of indices of rows of $C$ that are equal to $v^{t}_{k}$ and $I^{\prime}_{k}\subseteq[m_{1}]$ as the set of indices of rows of $C$ that are equal to $-v^{t}_{k}$. Now, for $k\in[2]$ we define $f^{\prime}_{k}:=\sum_{t\in I_{k}}f_{t}-\sum_{t\in I^{\prime}_{k}}f_{t}-\sum_{t\in I_{3}}f_{t}+\sum_{t\in I^{\prime}_{3}}f_{t}$. For $i\in\{3,\ldots,m_{2}+2\}$ we define $f^{\prime}_{i}:=f_{m_{1}+k-2}$. It is direct to check that this choice of $f^{\prime}$ satisfies the desired conditions.
Let $k=4$. Let $f=(f_{1},f_{2})$ with $f_{1}\in(\mathbb{Z}^{m_{1}})^{*}$ and $f_{2}\in(\mathbb{Z}^{m_{2}})^{*}$. For $j\in[n_{1}]$ the $f$-value of the $j$-th column sum of $M$ equals the sum of the $f_{1}$-value of the $j$-th column sum of $A$ and $f_{2}(v^{\prime})\cdot u_{j}$. For $(f_{1},f_{2}(v^{\prime}))\in(\mathbb{Z}^{m_{1}+1})^{*}$ we get that $\begin{pmatrix}A\\ u^{t}\end{pmatrix}$ is $w_{1}$-valued. Assume that two columns $j_{1},j_{2}$ of $\begin{pmatrix}A\\ u^{t}\end{pmatrix}$ would be equal to each other. Then also the $j_{1}$-th and the $j_{2}$-th column of $M$ would be equal to each other, a contradiction. Similarly, one proves the corresponding statement for $\begin{pmatrix}v^{t}\\ B\end{pmatrix}$.
3.4. Technical inequalities regarding the bound
We abbreviate $g(x):=\frac{(x+1)^{2}}{4}$. We define $h\,:\,\mathbb{Z}_{\geq 1}\to\mathbb{Z}_{\geq 1}$ as $h(x):=\lfloor g(x)\rfloor$ if $x\not=5$, and $h(5):=10$. Note that $h(5)=g(5)+1$, $h(x)=g(x)$ if $x\not=5$ is odd and $h(x)=g(x)-\frac{1}{4}$ if $x$ is even. We also remark that $h(i)<h(j)$ for $i<j$. Notice $h(1)=1$, $h(2)=2$, $h(3)=4$, $h(4)=6$, $h(5)=10$, $h(6)=12$, $h(7)=16$, $h(8)=20$, $h(9)=25$, $h(10)=30$. We will need the following auxiliary result.
Lemma 3.12.
  1. (1)
    Let $x,y\in\mathbb{Z}_{\geq 1}$. Then $h(x)+h(y)\leq h(x+y)$.
  2. (2)
    Let $x\in\mathbb{Z}_{\geq 2}$, $y\in\mathbb{Z}_{\geq 1}$. Then $h(x)+h(y+1)\leq h(x+y)$.
  3. (3)
    Let $x\in\mathbb{Z}_{\geq 3}$, $y\in\mathbb{Z}_{\geq 1}$. Then $h(x)+h(y+2)\leq h(x+y)$ except if $(x,y)$ is in
    $\{(3,1),(3,3),(5,1)\}.$
  4. (4)
    Let $x,y\in\mathbb{Z}_{\geq 2}$. Then $h(x+1)+h(y+1)\leq h(x+y)$ except if $(x,y)$ is in
    $\{(2,2),(2,4),(4,2)\}.$
Proof.
(1) The statement follows if $(g(x)+1)+(g(y)+1)\leq g(x+y)-\frac{1}{4}$, which is equivalent to $(x+y+1)^{2}-(x+1)^{2}-(y+1)^{2}\geq 9$, which holds [10] except if $y<\frac{5}{x}$. Hence, by symmetry we are left with checking $h(x)+h(y)\leq h(x+y)$ for $x,y\leq 4$.
(2) The statement follows if $(g(x)+1)+(g(y+1)+1)\leq g(x+y)-\frac{1}{4}$, which is equivalent to $(x+y+1)^{2}-(x+1)^{2}-(y+2)^{2}\geq 9$, which, as $x\geq 2$, holds [10] except if $y<\frac{13}{2(x-1)}$. Hence, we are left with checking $h(x)+h(y+1)\leq h(x+y)$ for $2\leq x\leq 7$ and $1\leq y\leq 6$.
(3) The statement follows if $(g(x)+1)+(g(y+2)+1)\leq g(x+y)-\frac{1}{4}$, which is equivalent to $(x+y+1)^{2}-(x+1)^{2}-(y+3)^{2}\geq 9$, which, as $x\geq 3$, holds [10] except if $y<\frac{9}{x-2}$. Hence, we are left with checking $h(x)+h(y+2)\leq h(x+y)$ for $3\leq x\leq 10$ and $1\leq y\leq 8$.
(4) The statement follows if $(g(x+1)+1)+(g(y+1)+1)\leq g(x+y)-\frac{1}{4}$, which is equivalent to $(x+y+1)^{2}-(x+2)^{2}-(y+2)^{2}\geq 9$, which, as $x,y\geq 2$, holds [10] except if $y<\frac{x+8}{x-1}$, which is at most $10$. Hence, by symmetry we are left with checking $h(x+1)+h(y+1)\leq h(x+y)$ for $2\leq x,y\leq 9$.
3.5. Proof of Theorem 1.4
Let $\mathbf{1}$ be the all ones vector. Note that $M$ is polytopal if and only if $M$ is $\mathbf{1}$-valued. We say $M$ is prepared if $M$ is polytopal and all columns are pairwise distinct.
Let $M$ be a prepared TU-matrix with $m$ rows. We will prove that the number of columns of $M$ cannot exceeds $h(m)$ by induction on the size.
From the TU-property it follows that one can multiply $M$ from the left by a unimodular transformation in $\text{GL}(\mathbb{Z},m)$ and permute the columns in a way such that the new matrix is of the form
$\begin{pmatrix}I_{r}&M^{\prime}\\ 0&0\end{pmatrix}$
where $(I_{r}|M^{\prime})$ is also a prepared TU-matrix. By induction hypothesis, we may assume that $M=(I_{m}|M^{\prime})$, where also $M^{\prime}$ is prepared (hence, by our assumption $M^{\prime}$ does not contain any standard basis vector as a column). In particular, $M$ has full rank $m$, and $M$ and $M^{\prime}$ have constant column sum $1$.
By the following remark we can assume that $m\geq 7$.
Remark 3.13.
By Proposition 2.9 any prepared TU-matrix with $m$ rows of rank $m$ defines an $(m-1)$-dimensional unimodular polytope. Hence, for $m\leq 6$ we can use Proposition 2.5 to determine the maximal number of columns and verify our statement.
We denote the number of columns of $M^{\prime}$ by $n$, so $M^{\prime}$ is an $m\times n$-matrix, $M$ is an $m\times(m+n)$-matrix and has size $2m+n$. We have to show $n\leq h(m)-m$.
Now, we make use of the following fundamental result (see [16] and [17]).
Theorem 3.14 (Seymour’s decomposition theorem).
Up to permutation of rows and columns, each TU-matrix is a network matrix, the transpose of a network matrix, a so-called sporadic matrix, or a $k$-sum for $k\in\{1,2,3,4\}$.
If $M^{\prime}$ is a network matrix or the transpose of a network matrix, it is easy to see that also $M$ is. Hence, we get our desired bound by applying Proposition 3.1 to $M$ or Theorem 3.2 to $M^{t}$. If $M^{\prime}$ is a sporadic matrix, see [17], then $n=5=10-5=h(m)-m$.
Hence, Seymour’s decomposition theorem applied to $M^{\prime}$ implies that after a permutation of rows and columns we may assume that $M^{\prime}$ is a $k$-sum for $k\in\{1,2,3,4\}$, where the $m_{1}\times n_{1}$-matrices $A$ and $m_{2}\times n_{2}$-matrices $B$ are given as in Subsection 3.3. Note that $m_{1}+m_{2}=m$ and $n_{1}+n_{2}=n$. We will consider these four cases separately.
Let $k=1$. As they are submatrices of smaller sizes, by induction hypothesis $\left(e_{1}\cdots e_{m_{1}}|A\right)$ and $\left(e_{m_{1}+1}\cdots e_{m}|B\right)$ are prepared TU-matrices such that $m_{1}+n_{1}\leq h(m_{1})$ and $m_{2}+n_{2}\leq h(m_{2})$. Hence, $n=n_{1}+n_{2}\leq h(m_{1})-m_{1}+h(m_{2})-m_{2}\leq h(m_{1}+m_{2})-(m_{1}+m_{2})=h(m_{1}+m_{2})-m$ by Lemma 3.12(1).
Let $k=2$. As $M^{\prime}$ is prepared and contains no standard basis vectors as column vectors, we have $m_{1}\geq 2$. Note that even $m_{1}\geq 3$ holds as no vector $\{-1,0,1\}^{2}$ has odd column sum except for $\pm$ a standard basis vector. Again, $(e_{1}\cdots e_{m_{1}}|A)$ is a prepared TU-matrix, hence, by induction hypothesis, $m_{1}+n_{1}\leq h(m_{1})$. By Lemma 3.11(1) applied to $M^{\prime}$, $\left(e_{2}\cdots e_{m_{2}+1}\begin{matrix}v^{t}\\ B\end{matrix}\right)$ is a prepared TU $(m_{2}+1)\times(m_{2}+n_{2})$-matrix. Here, note that $\begin{pmatrix}v^{t}\\ B\end{pmatrix}$ cannot contain a standard basis vector $e_{j}$ with $j\in\{2,\ldots,m_{2}+1\}$ as a column as otherwise by the definition of a $2$-sum the matrix $M^{\prime}$ would contain $e_{m_{1}+j-1}$ as a column. The induction hypothesis implies $m_{2}+n_{2}\leq h(m_{2}+1)$. Therefore, $n=n_{1}+n_{2}\leq h(m_{1})-m_{1}+h(m_{2}+1)-m_{2}\leq h(m_{1}+m_{2})-(m_{1}+m_{2})$ by Lemma 3.12(2).
Let $k=3$. As for $k=2$, we have $m_{1}\geq 3$ and $m_{1}+n_{1}\leq h(m_{1})$. As we otherwise would be in the case $k=1$, $C$ has a non-zero row $w^{t}$. If all non-zero rows of $C$ would be of the form ±$\pm w^{t}$, this would be the case $k=2$. Hence, we may assume that there are at least two non-zero and non-centrally-symmetric rows of $C$, say, of the form ±$\pm v^{t}_{1}$ and ±$\pm v^{t}_{2}$. By Lemma 3.11(2), $\left(e_{3}\cdots e_{m_{2}+2}\begin{array}[]{c}v^{t}_{1}\\ v^{t}_{2}\\ B\end{array}\right)$ is a prepared $(m_{2}+2)\times(m_{2}+n_{2})$ TU-matrix. Again, note that $\begin{pmatrix}v^{t}_{1}\\ v^{t}_{2}\\ B\end{pmatrix}$ cannot contain a standard basis vector $e_{j}$ with $j\in\{3,\ldots,m_{2}+2\}$ as a column as otherwise by the definition of a $3$-sum the matrix $M^{\prime}$ would contain $e_{m_{1}+j-2}$ as a column. Therefore, the induction hypothesis implies $m_{2}+n_{2}\leq h(m_{2}+2)$. Hence, $n=n_{1}+n_{2}\leq h(m_{1})-m_{1}+h(m_{2}+2)-m_{2}\leq h(m_{1}+m_{2})-(m_{1}+m_{2})$ by Lemma 3.12(3); note that all three exceptional cases cannot occur as they satisfy $m=m_{1}+m_{2}\leq 6$.
Let $k=4$. In the same way as above, $\begin{pmatrix}e_{1}\cdots e_{m_{1}}\begin{array}[]{c}A\\ u^{t}\end{array}\end{pmatrix}$ and $\begin{pmatrix}e_{2}\cdots e_{m_{2}+1}\begin{array}[]{c}v^{t}\\ B\end{array}\end{pmatrix}$ are prepared TU-matrices by Lemma 3.11(3). Hence, the induction hypothesis implies $n_{1}\leq h(m_{1}+1)-m_{1}$ and $n_{2}\leq h(m_{2}+1)-m_{2}$. Therefore, Lemma 3.12(4) implies $n=n_{1}+n_{2}\leq h(m_{1}+1)-m_{1}+h(m_{2}+1)-m_{2}\leq h(m_{1}+m_{2})-(m_{1}+m_{2})$ if $m_{1}\geq 2$ and $m_{2}\geq 2$.
Let us assume that $m_{1}=1$. If there would be an $i\in\{1,\ldots,n_{1}\}$ such that the $i$th coordinate of $u^{t}$ would be 00, then all coordinates except possibly the first one of the $i$th-column vector of $M^{\prime}$ would be 00, a contradiction to $M^{\prime}$ having column sums $1$ and not containing any standard basis vector as a column. Hence, every coordinate of $u^{t}$ is equal to ±$\pm 1$. Consider the $2\times n_{1}$ TU-matrix $\begin{pmatrix}A\\ u^{t}\end{pmatrix}$. If there are $i\not=j$ such that $A_{i}=A_{j}\not=0$, then because of the TU-property also $(u^{t})_{i}=(u^{t})_{j}$ holds, hence, the $i$th and the $j$th column of $M^{\prime}$ coincide, a contradiction. Therefore, if there are $i\not=j$ such that $A_{i}\not=0\not=A_{j}$, then $A_{i}=-A_{j}$, and by the same argument we also have $(u^{t})_{i}=-(u^{t})_{j}$, and thus, also the $i$th and the $j$th column of $M^{\prime}$ are centrally-symmetric, a contradiction to $M^{\prime}$ having column sums $1$. Also, we see in the same way that that there cannot exist $i\not=j$ such that $A_{i}=0=A_{j}$. This shows that $A$ has only at most two entries, i.e., $n_{1}\leq 2$. Hence, the size $m_{1}+n_{1}$ of $A$ is at most $3$, a contradiction to the assumption of Definition 3.9. This proves $m_{1}\geq 2$. In the same way, one shows $m_{2}\geq 2$. This finishes the proof.